OpenStudy (twopointinfinity):

Help please, will give medal. P(3, -2); csc 0 =

OpenStudy (anonymous):

I'm so sorry, I do not know xD

OpenStudy (anonymous):

csc means 1/sin right? so do you know how to find sin?

OpenStudy (anonymous):

if i am correct, then you would take your coordinates P(3, -2) and use it to find sin first. think of the coordinates like this:(cos, sin) so if you look at the graph, you see that you have to find the hypotenuse using the pythagorean theorem.

OpenStudy (anonymous):



OpenStudy (anonymous):

you wud plug in your variables in here a^2 + b^2 = c^2. (3)^2 + (-2)^2 = c^2 then solve. 9+4=c^2 13=c^2

OpenStudy (anonymous):